OpenStudy (anonymous):

Which circles have a circumference of 20 pi units? Choose all answers that are correct. A. a circle with a diameter of 40 units B. a circle with a radius of 10 units C. a circle with a diameter of 20 units D. a circle with a radius of 20 units

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions this is my last question

OpenStudy (anonymous):

I think something is missing from your question...

OpenStudy (anonymous):

like what

OpenStudy (anonymous):

Like in between pi and units

OpenStudy (anonymous):

Is there anything missing?

OpenStudy (anonymous):

no thats my bad to many spaces

OpenStudy (anonymous):

xd

OpenStudy (anonymous):

I think I know how to find the answer, but not so sure.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

oh so what do you think?

OpenStudy (welshfella):

circumference = 2*pi*r where r is radius

OpenStudy (welshfella):

so if radius = 10 then circumference = 2*pi*10 = ?

OpenStudy (welshfella):

also remember that the radius = half the diameter

OpenStudy (anonymous):

@welshfella What does it mean by 20 pi units?

OpenStudy (anonymous):

That's what I'm stuck on.

OpenStudy (welshfella):

2 * pi *10 = 20pi units

OpenStudy (welshfella):

20 pi = 20 times pi

OpenStudy (anonymous):

That's what I thought. :P

OpenStudy (welshfella):

yes the multiply sign is understood ( left out)

OpenStudy (anonymous):

so what are the answers @AnswerMyQuestions

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @welshfella 2 * pi *10 = 20pi units \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Hmm, @welshfella could you explain? I'm a little busy. :/

OpenStudy (anonymous):

the answers are b and c @AnswerMyQuestions

OpenStudy (anonymous):

They are?

OpenStudy (welshfella):

@naterboss - I've given you one answer already

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

2*pi*10 = 20pi

OpenStudy (anonymous):

thanks for the help and i got a 100

OpenStudy (welshfella):

now check to see if any of the other options have a radius of 10 - remembering that the diameter is 2 * radius

OpenStudy (welshfella):

yw